OpenStudy (anonymous):

PLEASE HELP :) Supplier A sells grain for $5.25 per bushel plus a delivery fee of $3 per bushel. Supplier B sells the grain for $6.50 per bushel plus a flat delivery fee of $1,500 each week. Which supplier would cost more if the production company bought 500 bushels of grain each week? A.Supplier B would cost $625 more than supplier B.Supplier B would cost $600 more than supplier C.Supplier A would cost $875 more than supplier D.Supplier A would cost $800 more than supplier

OpenStudy (anonymous):

first one would cost \[8.25\times 500\] second one would cost \[6.50\times 500+1500\] check which one is smallest

OpenStudy (anonymous):

i would use a calculator \[4125\] for the first one http://www.wolframalpha.com/input/?i=8.25*500

OpenStudy (anonymous):

THANK YOU SO MIUCH

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6.5*500%2B1500 4750 for the second yw

OpenStudy (anonymous):

then \[4750-4125=625\] go with A

OpenStudy (anonymous):

YEAH I WENT WITH A THANKS FOR YOUR HELP

OpenStudy (anonymous):

yw